1, Balance the following reaction in a basic solution:

I - + MnO42- → MnO2 + I2


2, Build a voltaic cell with one beaker containing potassium permanganate (KMnO4 ) solution with a platinum electrode and a second beaker containing zinc chloride with a zinc electrode. Determine:

a. The half reactions

b. Flow of electrons

c. Short-form notation

d. Electric potential (voltage)


1
Expert's answer
2021-01-11T03:47:23-0500


1- 4H2O + 6I- +2MnO4- = 3I2 + 8OH- +2MnO4


2- a)- at anode oxidation reaction-

Zn(S) = Zn2+ + 2e-

At cathode Reduction reaction-

Nl+2(aq) +2e_ = Ni(S)

b)- From anode to cathode

c)- Zn(s) I ZnCl2(aq) II Ni(No3)2(aq) Ni(S)

d)-Electric Potential =

"E= E^o - \frac{RT}{2F}ln \frac{[Zn^+2][Ni]}{[Zn][Ni^+2]}"